The current density in a cylindrical wire of radiu
Answer #1

Here,

R = 2 mm

J = 2 *10^5 A/m^2

for the current through wire between R/2 and R = J * area

the current through wire between R/2 and R = J * pi (R^2 - (R/2)^2)

the current through wire between R/2 and R = 2 *10^5 * pi * (0.002^2 - (0.002/2)^2)

the current through wire between R/2 and R = 1.885 A

the the current through wire between R/2 and R is 1.885 A

-----------

b)

J = a r^2

J = 3 *10^11 * r^2

For the current

I = integration(J * 2pir * dr)

I = integration(3 *10^11 * 2 * pi * r^3 . dr)

I = 3 *10^11 * 2pi * [r^4/4]

I = 3 *10^11 * 2pi * [0.002^4 - 0.001^4]/4

I = 7.07 A

the current in the loop is 7.07 A